# supplementary

### v1

@@ 3.4 Spurious Feature Suppression
Spurious feature suppression during inference is a two step process. Given a test image we first use the learned prototypical patches to identify the parts of the image containing spuriously correlated attributes. Then we use simple token ablation to eliminate these patches from the image, ensuring the model does not rely on spurious attributes.
@@ 3.4.1 Spurious Feature Identification during Inference
Detecting which patches - if any - in a test image contain spuriously correlated features can be achieved using a supervised classification approach. While simple approaches that rely on similarity in the feature space to the previously computed prototypical patches worked well, we found that a K-Nearest Neighbors (KNN) classifier in the key space representation proved superior in both performance and efficiency. Simply put, we construct a balanced training set using a bank of key vectors from high $P(p)$ scoring patches as positive samples (class $1$) from the selected shortcut cluster, combined with an equal number of low $P(p)$ scoring key vectors from the non-shortcut cluster as negative samples (class 00). The same key embeddings are computed for all patches in the test-image, and the trained KNN classifier determines which patches to ablate (see Figure 1 H).
@@ 3.4.2 Spurious Feature Ablation
Our framework takes advantage of the fact that the transformer architecture accepts an arbitrary number of patches by removing any patch that the previous step detected as containing a spuriously correlated features from the tokenized representation. This ablation is done after the positional embedding, therefore the embedding of all other image patches remain unaffected. While more sophisticated approaches are viable, patch ablation is likely the simplest and most computationally efficient approach. As discussed in Section 5 it also proved to be highly effective as images that do not contain spurious attributes are left largely unaffected (see Table 2), therefore unlike in previous research, we avoid any adverse effects on accuracy of groups that do not contain the spurious attribute.
As demonstrated in figure 1, the proposed approach is highly modular, consisting of a series of independent components. While multiple alternatives of each component achieved comparable results, this setup was ultimately chosen due computational efficiency and simplicity.
3.5 Deep Feature Reweighting
We also extend our framework by retraining the last transformer layer using the ablated images from the last step of our approach. This last layer retraining is a analogous to the popular Deep Feature Reweighting approach for shortcut mitigation in convolutional neural networks [16]. Despite DFR being a powerful approach, researchers have found evidence that it is not sufficient to completely mitigate shortcut learning [18], instead arguing that only by retraining with data from which the spurious features were eliminated one ensures that the model classifier head effectively forgets the spurious attributes completely [24]. While editing training data is complex and computationally expensive, utilizing the ablated images from the last step of our approach (and leveraging the fact that transformers do not require a fixed number of input tokens) we can effectively perform the same retraining for minor additional cost.
In conclusion, integrating DFR into our framework eliminates many of it’s drawbacks, resulting in significant performance improvements for a relatively small computational cost (see Table 1).
@@ 4 Experimental Setup
We benchmark our framework (with and without DFR) against two popular shortcut mitigation methods; Deep Feature Reweighting [16] and Just Train Twice [20], which require group annotations. The Optuna framework was used for hyperparameter optimization.
@@ 4.1 Models
While the shortcut learning literature focuses mostly on popular convolutional neural networks, Transformer based architectures have rapidly surpassed them as state-of-the-art in computer vision. Preliminary work into shortcut learning in visual transformers have found that they are relatively robust to spurious correlations [15]. Our work focuses on the ViT B-16 visual transformer architecture which remains ubiquitous and competitive in computer vision research [14, 15]. The baseline model was pretrained using the common ILSVRC subset of ImageNet [^2] Pretrained weights were downloaded from the TorchVision library, and fine-tuning was done on each binary classification dataset individually.
@@ 4.2 Datasets
We benchmark our approach on three different datasets. Similarly, to previous research [15] we found that vision transformers are not susceptible to many of the shortcuts found in popular datasets (for instance, worst group accuracy in the popular waterbird dataset was over ${}90\%$). Therefore, we utalized newer, more challenging, datasets. Moreover, considering interpretability and computational accessability are especially important in medical diagnostic tasks, we focus on medical datasets. Each experiment was repeated with three sequential seeds.
@@ 4.2.1 ISIC
The International Skin Imaging Collaboration (ISIC) skin lesion dataset [10] is a popular dataset in the shortcut learning literature [18, 24]. This dataset contains images of malignant and benign tumors (2000 training samples in each class). Of the benign samples, 1000 images contain colored bandages - while only 10 (artificially generated) are to be found in the malignant training samples - thereby constituting a spuriously correlated attribute. The validation set is 10% of the size of the train set with a similar group distribution. Additionally we generated 100 malignant tumor images which we introduced bandages into (see Appendix 1). The worst group accuracy for this dataset refers to the ratio of correctly classified malignant samples with bandage.
@@ 4.2.2 Knee Radiographs
The knee Osteoarthritis radiograph dataset [7] contains knee joint x-ray images of 4796 participants suffering from various degrees of Osteoarthritis. Due to ViTs outperforming CNNs, unlike in previous work [1] we used the more challenging task of ‘no’ and ‘moderate’ Osteoarthritis classification. Following previous work [1, 11] we added a hospital tag to 50% and 2.5% of the images from healthy and arthritic patients respectively. The training and validation sets contained 1000 and 200 samples per class respectively (with the same group distribution). For testing we generated 100 samples of each class with and without the spurious marker. The worst group in this dataset refers to moderate arthritis images with the added radiographic marker.
@@ 4.2.3 ImageNet-W
Zhiheng et al. [19] discovered that the carton class in the popular ImageNet dataset [13] is spuriously correlated with the presence of Chinese characters due to watermarks. We utilize this already existing shortcut in combination with the visually similar storage image-net class which does not contain any watermarks. We used 300 of the spuriously correlated carton class training images and a similar amount of the storage class images for training and 100 images from each for validation. For testing we used 100 images without watermarks of both the storage and carton class, and the same images with added watermarks (using code provided by Zhiheng et al.). The group with the worst accuracy was the artificially created watermarked storage samples.
5 Results
@@
SCM 75.0 $\pm$ 2.4 93.1 $\pm$ 1.0 77.5 $\pm$ 2
+ DFR 87.0 $\pm$ 6.5 95.3 $\pm$ 1.0 88.2 $\pm$ 3
Table 1: Worst and average group accuracy (mean and standard deviation) after shortcut mitigation. The baseline is a pretrained Vit-B-16 fine-tuned on each dataset. For shortcut mitigation we use Deep Feature Reweighting (DFR), Just Train Twice (JTT) and our shortcut mitigation (SCM) approach with and without DFR. We also report total runtime seconds (including inference).
Our unsupervised ablation based shortcut mitigation framework (SCM) is competitive with methods that require group annotation both in terms of worst and average group accuracy. Moreover, integrating deep feature reweighting further increased performance across all datasets.
@@ 5.1 Computational Efficiency
We designed our framework with computational efficiency in mind. We ran the unsupervised versions of all experiments on a standard 16GB RAM Apple MacBook Pro (M3). As can be observed in table 1, our approach provides significant boost to worst group and average accuracy in a matter of minutes. Hence, our approach is especially promising for end-users looking to interactively explore their datasets and models to identify and eliminate shortcut learning.
@@ 5.2 Spurious Feature Ablation
After identifying prototypical patches that encode spurious attributes embedding similarity (in the key space) is used to detect and ablate patches containing spurious attributes during inference. As can be observed in Table 2, images that do not contain spurious attributes were mostly unaffected by our method. Examples of the prototypical patches encoding spurious attributes in each dataset can be found in Figure 2, we also manually validated that the spurious attributes were always correctly identified.
@@
Table 2: Percentage (mean $\pm$ std) of samples where we mask at least one token in the groups where i) the shortcut is present (SP) and ii) the shortcut is not present (NS).
@@
Table 3: Concept descriptions for shortcut and non-shortcut clusters generated by different MMLLMs. Human experts can easily identify shortcut features based on the generated text. For more information regarding the prompts and setup see appendix 5
@@ 5.3 Spurious Attribute Detection
For a shortcut detection framework to be effective, it is necessary for it to capture attribute concepts that are semantically meaningful. Indeed, previous research found that most shortcut detection methods fail to provide users with sufficient information to identify the spurious attributes [1]. Motivated by previous research we conducted an analogous study. We use the patch descriptions generated by the multi-modal LLaVA model [21], and then further refined the results using another text only LLM.
User Study: For each dataset, we used LLaVA to generate descriptions given the prototypical patches of the cluster identified to most likely represent a shortcut. Three LLMs were then prompted to identify potential shortcuts based on the LLaVA outputs. We manually verified that the correct attribute is always present in the top three responses for each LLM. Participants were given a short description of the datasets and asked to identify which response most likely describes a shortcut (for additional details see Appendix 5).
Results: 34 participants took part in the study. Each was given the top three responses from LLaMA-70B, LLaMA-8B, and Mixtral8x7B and descriptions of the datasets (with no visual exemplars). On average participants identified the spurious attributes with accuracy significantly above chance ($\mu=86.8\%$, $SD=17.6\%$, chance level $=51.3\%$, one-tailed t-test statistics $t(33)=1670.7647$, $p<0.0001$), demonstrating that the responses generated by the LLMs were indeed helpful for the end-user even without prior familiarity with the dataset, or even visual examples of the patch prototypes. Considering that domain experts using our framework will also have access to the visual prototypical samples, this accuracy can be perceived as a lower bound. Overall, these results indicate that explainable AI techniques that employ LLMs for concept guidance [31] can be successfully extended for shortcut detection.
@@ 6 Discussion
As demonstrated in Section 5, our approach offers an efficient framework for detecting and mitigating shortcuts in transformers. Beyond performance, several design choices distinguish our method from recent shortcut learning research: As detailed in Section 2, previous unsupervised methods often focuses on mitigating spurious shortcuts without explicitly identifying them [29, 33, 2]. In contrast, our design ensures that spurious attribute detection is explicit, and can readily incorporate human input. This is achieved through the integration of prototype learning [6, 23] and MLLM-guided concept identification [5].
This is important for multiple reasons: First, surveys demonstrate that domain experts are reluctant to trust black-box models, regardless of accuracy [8]. Second, the identification of shortcuts is task-dependent; for example, Chinese characters may be spurious attributes in object classification but core features in geographic location prediction. We hope this work further advances the role of human interpretability in shortcut learning research.
Despite the encouraging results, our approach has several limitations which we look forward to address in future work. First and foremost, recent research has demonstrated that spurious correlations come in droves and eliminating one enhances the classifier’s reliance on other, previously hidden attributes [19]. We have encountered a similar situation when applying our approach to the CelebA dataset. Instead of the prototypical patches capturing gender, they contained concepts such as cloth color, jewelry, and the presence of neckties (see Appendix 3). While medical data is often acquired in controlled environments and is relatively clean, future research focusing on shortcut learning in complex datasets is necessary.
7 Conclusion

### v2

@@ 3.4 Spurious Feature Suppression
Spurious feature suppression during inference is a two-step process. Given a test image, we first use the learned prototypical patches to identify the parts of the image that contain spuriously correlated attributes. Then we use simple token ablation to eliminate these patches from the image, ensuring that the model does not rely on spurious attributes.
@@ 3.4.1 Spurious Feature Identification during Inference
Detecting which patches, if any, in a test image contain spuriously correlated features is achieved using a supervised classification approach. Although simple approaches based on feature-space similarity to the previously computed prototypical patches worked well, a K-Nearest Neighbors (KNN) classifier in the key-space representation had superior performance and efficiency. Simply put, we construct a balanced training set using a bank of key vectors from high $P(p)$ scoring patches as positive samples (class $1$) from the selected shortcut cluster, combined with an equal number of low $P(p)$ scoring key vectors from the non-shortcut cluster as negative samples (class 00). Key embeddings are computed for all test image patches, and the trained KNN classifier determines which patches to ablate (Figure 1 H).
@@ 3.4.2 Spurious Feature Ablation
Our framework takes advantage of the fact that the transformer architecture accepts an arbitrary number of patches by removing any patch that the previous step detected as containing spuriously correlated features from the tokenized representation. This ablation is done after the positional embedding; therefore, the embedding of all other image patches remains unaffected. Although more sophisticated approaches are viable, patch ablation is likely the simplest and most computationally efficient approach. As discussed in Section 5, it also proved to be highly effective as images that do not contain spurious attributes are left largely unaffected (see Table 2); therefore, unlike previous research, we avoid any adverse effects on accuracy of groups that do not contain the spurious attribute.
3.4.3 Last Layer Retraining
Having eliminated image patches that contain spurious features, the last step involves retraining to increase the models reliance on core features. We found that retraining the last dense transformer layer using the embedding of the images post-ablation achieved improved performance for little computational cost. Note that this last step is not analogous to the popular Deep Feature Reweighting approach for shortcut mitigation in convolutional neural networks, as we do not rely on group annotation [17]. Moreover, despite DFR being a powerful approach, researchers have found evidence that it is not sufficient to completely mitigate shortcut learning [19], instead arguing that only retraining with data from which spurious features were eliminated ensures that the model classifier head effectively forgets the spurious attributes completely [25]. While editing training data is complex and computationally expensive, by utilizing the ablated images from the previous step of our approach (and leveraging the fact that transformers do not require a fixed number of input tokens), we can effectively ensure that the retrained model relies only on core features.
@@ 4 Experimental Setup
We benchmark our framework against popular shortcut mitigation methods that require group annotations: subsampling groups (SUBG) [16], Deep Feature Reweighting [17], and Just Train Twice [21]. Following standard practice, we use worst and average group accuracy for evaluation (where groups are defined by label and shortcut presence). The Optuna framework was used for hyperparameter optimization.
@@ 4.1 Models
While the shortcut learning literature focuses mostly on popular convolutional neural networks, Transformer-based architectures have rapidly surpassed them as state-of-the-art in computer vision. Our work focuses on the ViT B-16 vision transformer architecture which remains ubiquitous and competitive in computer vision research [14, 15]. The baseline model was pretrained using the common ILSVRC subset of ImageNet [^2] Pretrained weights were downloaded from the TorchVision library, and fine-tuning was done on each binary classification dataset individually.
@@ 4.2 Datasets
We benchmark our approach on three different datasets. Similarly, to previous research [15] we found that vision transformers are not susceptible to many of the shortcuts found in popular datasets (for instance, worst group accuracy in the popular waterbird dataset was over ${}90\%$, see Appendix C). Therefore, we utilized newer, more challenging, datasets. Moreover, considering interpretability and computational accessibility are especially important in medical diagnostic tasks, we focus on medical datasets. Each experiment was repeated with three sequential seeds.
@@ 4.2.1 ISIC
The International Skin Imaging Collaboration (ISIC) skin lesion dataset [10] is a popular dataset in the shortcut learning literature [19, 25]. This dataset contains images of malignant and benign tumors (2000 training samples in each class). Of the benign samples, 1000 images contain colored bandages, whereas only 10 (artificially generated) are to be found in the malignant training samples, thus constituting a spuriously correlated attribute. The validation set is 10% of the size of the train set with a similar group distribution. In addition, we generated 100 malignant tumor images in which we inserted bandages (see Appendix A). The worst group accuracy for this dataset refers to the ratio of correctly classified malignant samples with bandage.
@@ 4.2.2 Knee Radiographs
The knee osteoarthritis radiograph dataset [7] contains knee joint X-ray images of 4796 participants suffering from various degrees of osteoarthritis. Due to ViT outperforming CNNs, unlike in previous work [1] we used the more challenging task of classification of ’no’ vs. ’moderate’ osteoarthritis. Following previous work [1, 11] we added a hospital tag to 50% and 2.5% of the images from healthy and arthritic patients respectively. The training and validation sets contained 1000 and 200 samples per class, respectively (with the same group distribution). For testing, we generated 100 samples of each class with and without a spurious marker. The worst group in this dataset refers to moderate arthritis images with the added radiographic marker.
@@ 4.2.3 ImageNet-W
Zhiheng et al. [20] discovered that the carton class in the popular ImageNet dataset [13] is spuriously correlated with the presence of Chinese characters due to watermarks. We utilize this already existing shortcut in combination with the visually similar storage image-net class, which does not contain any watermarks. We used 300 of the spuriously correlated carton class training images and a similar amount of the storage class images for training and 100 images from each for validation. For testing we used 100 images without watermarks of both the storage and carton class, and the same images with added watermarks (using code provided by Zhiheng et al.). The group with the worst accuracy was the artificially created watermarked storage samples.
@@
ASM (ours) 87.0 $\pm$ 6.5 95.3 $\pm$ 1.0 88.2 $\pm$ 3
w/o retraining 75.0 $\pm$ 2.4 93.1 $\pm$ 1.0 77.5 $\pm$ 2
Table 1: The baseline is a pretrained Vit-B-16 fine-tuned on each dataset. Our ablation-based shortcut mitigation (ASM) approach (with and without last layer retraining) is benchmarked against subsampling groups (SUBG), Deep Feature Reweighting (DFR), Just Train Twice (JTT). We report Worst and average group accuracy (mean $\pm$ std) and total runtime seconds (including inference).
5 Results
Our unsupervised ablation-based shortcut mitigation framework (ASM) outperforms popular methods that require group annotation both in terms of worst- and average-group accuracy. Moreover, an ablation experiment demonstrates that simply eliminating patches containing spurious features, without any retraining, already results in a significant performance increase.
@@ 5.1 Computational Efficiency
We designed our framework with computational efficiency in mind. We ran the unsupervised versions of all experiments on a standard 16GB RAM Apple MacBook Pro (M3). As can be observed in Table 1, our approach provides significant boost to worst group and average accuracy in a matter of minutes. Hence, our approach is especially promising for end users looking to interactively explore their datasets and models to identify and eliminate shortcut learning.
@@ 5.2 Spurious Feature Ablation
Examples of prototypical patches identified as encoding spurious attributes in each dataset can be found in Figure 2. Embedding similarity (in the key space) to these prototypical patches is used to detect and ablate patches containing spurious attributes. As can be observed in Table 2, images that do not contain spurious attributes were mostly unaffected by our method. These results, in conjunction with the ablation study results (where only patch ablation without retraining was used), indicate that our detection component correctly identifies spurious attributes in the stimuli.
@@
Table 2: Percentage (mean $\pm$ std) of samples where we mask at least one token in the images in which the shortcut is present (SP) or no shortcut is present (NS).
@@
Table 3: Concept descriptions for shortcut and non-shortcut clusters generated by different MMLLMs. Human experts can easily identify shortcut features based on the generated text. For more information regarding the prompts and setup see appendix E
@@ 5.3 Spurious Attribute Detection
For a shortcut detection framework to be effective, it is necessary to capture attribute concepts that are semantically meaningful. In fact, previous research found that most shortcut detection methods do not provide users with enough information to identify spurious attributes [1]. Motivated by previous research, we conducted an analogous study. We used the patch descriptions generated by the multimodal LLaVA model [22], and then further refined the results using a text-only LLM.
User Study: For each dataset, we used LLaVA to generate descriptions given the prototypical patches of the cluster identified to most likely represent a shortcut. Then three LLMs were prompted to identify potential shortcuts based on the LLaVA outputs. We manually verified that the correct attribute is always present in the top three responses for each LLM. Participants were given a description of the task and asked to identify which response most likely describes a shortcut (for additional details see Appendix E).
Results: 50 participants took part in the study. Each was given the top three responses from LLaMA-70B, LLaMA-8B, and Mixtral8x7B and descriptions of the datasets (with no visual exemplars). On average participants identified the spurious attributes with accuracy significantly above chance ($\mu=91.2\%$, $SD=18.4\%$, chance level $=51.3\%$, one-tailed t-test statistics $t(49)=15.33$, $p<0.0001$), demonstrating that the responses generated by the LLMs were indeed helpful for the end-user even without prior familiarity with the dataset, or even visual examples of the patch prototypes. Considering that domain experts using our framework will also have access to the visual prototypical samples, this accuracy can be perceived as a lower bound. Overall, these results indicate that explainable AI techniques that employ LLMs for concept guidance [32] can be successfully extended for shortcut detection.
@@ 6 Discussion
As demonstrated in Section 5, our approach offers an efficient framework for detecting and mitigating shortcuts in transformers. Beyond performance, several design choices distinguish our method from recent research on shortcut learning: As detailed in Section 2, previous unsupervised methods often focus on mitigating spurious shortcuts without explicitly identifying them [30, 34, 2]. In contrast, our design ensures that spurious attribute detection is explicit and can readily incorporate human input. This is achieved through the integration of prototype learning [6, 24] and MLLM-guided concept identification [5].
This is important for multiple reasons: First, surveys demonstrate that domain experts are reluctant to trust black-box models, regardless of model accuracy [8]. Second, the identification of shortcuts is task dependent, making human supervision crucial; for example, Chinese characters may be spurious attributes in object classification but core features in geographic location prediction. We hope that our work inspires further research that focuses on the role of human interpretability in the mitigation of shortcut learning.
Finally, as noted in previous research [15], we found that transformers are resistant to many shortcuts that affect convolutional networks, such as in CelebA and Waterbirds (Appendix C). Hence, our work is a needed step toward developing techniques and benchmarks that address the unique characteristics of shortcut learning in transformers.
7 Conclusion and Future Work
@@
Despite encouraging results, our approach has several limitations which we look forward to addressing in future work. Specifically, recent research has demonstrated that eliminating one spurious attribute could enhance the classifier’s reliance on other, previously hidden attributes [20]. We have encountered a similar situation when applying our approach to the CelebA dataset: Instead of the prototypical patches capturing gender, they focused on cloth color, jewelry, and the presence of neckties (see Appendix C). Although medical data are often acquired in controlled environments and are relatively clean, future research focusing on shortcut learning in complex datasets is necessary.
Acknowledgments
This work was funded by the Hessian.AI Connectom Networking and Innovation Fund 2024 and the European Union (ERC, TAIPO, 101088594 to F.B.) grant. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or ERC. Neither the European Union nor the granting authority can be held responsible for them.
Supplementary Material: Efficient Unsupervised Shortcut Learning Detection and Mitigation in Transformers
Appendix A ISIC Dataset
To assess the classifier’s performance on ISIC images of malignant tumors with colored bandages (representing the worst group performance), we manually added colored bandages to malignant tumor images from the validation set. This was done by cutting patches from unused training images using the background removal model tracer_b7, available as an API on Replicate. We obtained cutouts of colored patches, which were then layered onto malignant tumor samples using GIMP, varying the size, color, and location of the patches based on the training distribution.
Appendix B Knee Radiographs
Radiographic markers are frequently used to indicate the orientation and body part of the image. We obtained a cutout of an R (right body part) and L (left body part) marker from a hand x-ray image which we cutout with GIMP. We then automatically inserted the marker based on which knee (left or right) is visible in the image and varied in which corner (upper left and right as well as lower left and right) the marker is being added. We also added some slight rotation (between -5 and 5 degrees) to the added marker to introduce a more natural shortcut.
This follows the methodology introduced by Adebayo et al. [1] where they added a text (”MGH”) as an artifical hospital tag on the image. Our approach occurs frequently in a variety of datasets which makes it even more natural.
Appendix C Commonly used datasets
Recent research suggests that Vision Transformers are quite robust against spurious correlations in commonly used datasets such as Waterbirds and CelebA. Ghosal et al. [15] finetuned a ViT B-16 on Waterbirds resulting in a 96.75% average group accuracy and a 89.3% worst group accuracy. Similarly, they finetuned a ViT B-16 on CelebA resulting in an average group accuracy of 97.4% and a worst group accuracy of 94.1%.
We replicated their CelebA results by fine-tuning a ViT B-16 with the same hyperparameters as for all our other datasets, achieving above 90% AGA and WGA. Running our shortcut detection and mitigation framework on this dataset also replicates Li et al.’s findings [20] that eliminating one shortcut in this dataset will result in another being chosen by the model.
We detected a multitude of shortcuts which we were able to confirm as actual spurious correlations via the obtained labels (CelebA contains 40 labeled features including ”necktie”, ”glasses” and ”heavy-makeup”). A sample of a detected cluster can be seen in Figure 3.
We again ran LLaMa-70b to confirm the shortcuts in the prototypical patches and obtained the following summaries:
  • ”The model seems to focus on blonde hair, long hair, ponytails, and beards/mustaches, which might be shortcuts for identifying women or men.”
  • ”The model appears to focus on facial features like smiles, eyeshadow, lipstick, and moles, as well as accessories like glasses, ties, and shirts”
Figure 3: One of the many detected spurious correlations in the CelebA dataset are shirt collars for men.
Appendix D Clustering
We calculated the average cluster accuracies for all datasets for all three sequential seeds (see Table 4). As described in Section 3.2.1, clustering results are refined using a patch similarity measure, ensuring robustness even if the initial clustering is imperfect. Although we lack ground-truth annotations for shortcut feature locations, manual inspection of 200 ISIC prototypical patches confirms that all patches in the shortcut cluster contain the expected spurious feature.
Accuracy (%) $\uparrow$
ISIC 76.0
Knee Radiographs 100.0
ImageNet-W 91.6
Table 4: Average KNN clustering accuracy for all three sequential seeds, using two clusters.
We also couldn’t see any improvements in overclustering, as proposed by Sohoni et al. [29] (see Table 5).
[…]
2 Clusters 61.0 $\pm$ 2.4 87.3 $\pm$ 1.2
3 Clusters 56.7 $\pm$ 3.4 86.2 $\pm$ 1.7
Table 5: Worst and average group accuracy (mean and standard deviation) after shortcut mitigation with different clusters.
Appendix E User Study
We used the Replicate API to easily obtain results for multiple open-source LLMs. We decided on using three recently released open-source models with different parameter sizes:
  • LLaMa3-8b: The smallest open source LLaMa model with 8 billion parameters.
  • Mixtral8x7b: Mixture of experts architecture with 13 billion parameters.
  • LLaMa3-70b: The LLaMa model with 70 billion parameters.
We prompted all three models with the same prompt: ”I extracted patches from images in my dataset where my model seems to focus on the most. I let an LLM caption these images for you. I am searching for potential shortcuts in the dataset. Can you identify one or more possible shortcuts in this dataset? Describe it in one sentence (only!) and pick the most significant. No other explanations are needed. Descriptions:” followed by the captions that we obtained via the LLaVa-13b model. The LlaVa-13b model was prompted with the prototypical patches and the text prompt ”What is in this picture? Describe in a few words.”.
The study was conducted using a google forms. The participants were prompted with the dataset description and asked to identify which of the three responses was likely describes a spuriously correlated attribute. Often there were multiple correct answers, hence chance performance was $51.3\%$. Note that we used responses only based on the cluster our unsupervised method identified as the one most likely to contain spurious correlations. Therefore, the results of the survey validate that LLMs are capable of generating concepts that distill the properties captured by the patch prototypes.